# An Optimal Design Framework for Lasso Sign Recovery

### v2

@@ 3 Optimal Correlation Matrices
The probabilities for events (7) and (8) for $\textbf{C}\in\mathcal{C}$ are given in the following lemma:
@@
$\boldsymbol{u}$ $\sim N\left(\frac{\lambda\sqrt{n}}{1-c}\left[\boldsymbol{1}-z_{\mathcal{A}}\gamma\boldsymbol{z}_{\mathcal{A}}\right],\ \frac{1}{1-c}\left[\textbf{I}_{k}-\gamma\boldsymbol{z}_{\mathcal{A}}\boldsymbol{z}_{\mathcal{A}}^{T}\right]\right)$
$\boldsymbol{v}$ $\sim N\left(\lambda\sqrt{n}z_{\mathcal{A}}\gamma\boldsymbol{1},(1-c)\left[\textbf{I}+\gamma\textbf{J}\right]\right)$
@@
The criterion $\Phi_{\lambda}(\textbf{X}\,|\,k,\beta)$ meets the latter condition of Lemma 2 and assumes $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$, leading to the random vectors
$\boldsymbol{u}\sim N\left(\frac{\lambda\sqrt{n}}{1+c(k-1)}\boldsymbol{1},\frac{1}{1-c}\left[\textbf{I}-\gamma\textbf{J}\right]\right),\qquad\boldsymbol{v}\sim N\left(\lambda\sqrt{n}k\gamma\boldsymbol{1},(1-c)\left[\textbf{I}+\gamma\textbf{J}\right]\right)\ .$
The corresponding criterion defined with respect to $\mathcal{C}$ and for a fixed $\lambda$ is denoted
$\psi_{\lambda}(c\,|\,k,\beta)=P(S_{\lambda}\,|\,c,\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1})\times P(I_{\lambda}\,|\,c,\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1})\ .$ (16)
@@
$\psi_{\lambda}^{\pm}(c\,|\,k,\beta)=\frac{1}{2^{k-1}}\sum_{\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}^{\pm}}P(S_{\lambda}\,|\,c,\boldsymbol{\beta}_{\mathcal{A}}=\beta\tilde{\boldsymbol{z}})\times P(I_{\lambda}\,|\,c,\boldsymbol{z}_{\mathcal{A}}=\tilde{\boldsymbol{z}})\ .$ (17)
@@
The criteria $\psi_{\lambda}$ and $\psi_{\lambda}^{\pm}$ involve a single variable, but are still challenging to optimize analytically. Numerical optimization of these criteria, however, is straightforward and computationally efficient. We demonstrate numerical optimization of these new criteria for the situation of $p=10$, $k=4$, $\beta=2$, and $n=10$. Figure 2 shows the contour plots of $\psi_{\lambda}$ and $\psi_{\lambda}^{\pm}$. The optimal $c$ values for the corresponding $\psi_{\Lambda}$ and $\psi_{\Lambda}^{\pm}$ are $0.14$ and 00, respectively. The resulting optimal C matrices then match the ideal forms for the $Var(s+)$ and $UE(s^{2})$-criterion. While these ideal C are not possible for SSDs, this example provides further justification for these two heuristic SSD criteria for sign recovery under the lasso.
Figure 2: Contour plots of sign recovery probabilities when C is completely symmetric across possible values of the off-diagonal $c$ and $\log(\lambda)$ where $n=10$ and $k=4$ factors are active with $\beta=2$.
@@
We now study the behavior of $\psi_{\lambda}$ and $\psi^{\pm}_{\lambda}$ in a neighborhood about $c=0$. Direct application of the multivariate Leibniz integral rule gives the following lemma.
@@
For all $\mathcal{A}$ where $|\mathcal{A}|=k$ and $\beta>0$, $\frac{d}{dc}P(I_{\lambda}|c,\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1})\big{|}_{c=0}=0$ for all $\lambda$ and $\frac{d}{dc}P(S_{\lambda}|c,\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1})\big{|}_{c=0}\geq 0$ for $\lambda$ satisfying
$2\lambda_{n}$ $\geq\frac{g(\tau_{n})}{G(\tau_{n})}\ ,$ (18)
where $G(\cdot)$ and $g(\cdot)$ represent the $N(0,1)$ cumulative distribution function and probability density function, respectively, with $\lambda_{n}=\lambda\sqrt{n}$ and $\tau_{n}=\sqrt{n}(\beta-\lambda)$.
A direct consequence of Lemma 3 is that $\psi_{\lambda}(0|k,\beta)$ can be improved upon by some $c>0$. We state this important result as a theorem.
@@
Applying this result to the situation shown in Figure 2 with known sign, inequality (18) holds for $\log(\lambda)\geq-22.763$, which covers the entire region of $\lambda$ values for which $\psi_{\lambda}(c)>0$. It follows then that some $c>0$ will maximize the summary measures $\psi_{\Lambda}(c)$ or $\psi_{\max}(c)$. Theorem 2 thus provides a mathematically rigorous justification that the ideal designs under the $Var(s+)$-criterion maximize the probability of sign recovery for known sign.
@@
$\begin{split}\frac{q}{\binom{k}{2}}\ \frac{\lambda_{n}g(\lambda_{n})}{G(\Delta_{n})}&\left(k(1-\lambda_{n}^{2})+(q-1)\frac{\lambda_{n}g(\lambda_{n})}{G(\Delta_{n})}\right)\leq\\ &\frac{g(\tau_{n})}{G(\tau_{n})}\left(\beta_{n}+\lambda_{n}+\lambda_{n}^{2}\tau_{n}-\left[\frac{\beta_{n}^{2}-\lambda_{n}^{2}}{2}+\beta_{n}\lambda_{n}\right]\frac{g(\tau_{n})}{G(\tau_{n})}\right)\end{split}$ (19)
where $\beta_{n}=\beta\sqrt{n}$ and $G(\Delta_{n})=G(\lambda_{n})-G(-\lambda_{n})$.
Applying this result to the situation shown in Figure 2 with unknown sign, inequality (69) holds for $\log(\lambda)\in[-0.988,0.640]$. For $\log(\lambda)$ outside this region, there are clearly some $c>0$ for which $\psi_{\lambda}^{\pm}(c)>0$, although the probabilities are relatively small. These small probabilities do not influence $\psi_{\max}^{\pm}$ and have minimal influence on $\psi_{\Lambda}^{\pm}$, making $c=0$ a global maximum for both criteria. We conjecture that generally $c=0$ and some $c>0$ are global maxima for the unknown and known sign criteria, respectively.
@@ 4 Fast Design Construction with HILS
[…]
Our proposed construction algorithm first identifies a set of $ncand$ candidate designs through the Pareto front optimization search algorithm from Cao et al. (2017) based on our two lasso-motivated criteria. We then employ our exact lasso sign recovery criteria from Section 2 as a secondary criterion, as it more accurately measures a design’s screening capabilities, to rank the $ncand$ designs and choose the best one. We refer to this algorithm as the Heuristic-Initiated Lasso Sieve, or HILS.
@@
Although HILS can significantly reduce the number of $\Phi_{\lambda}$ evaluations compared to constructing an optimal design using, for instance, a coordinate exchange algorithm, there are settings when even a more modest number of evaluations is not computationally feasible. The same issue was encountered by Singh and Stufken (2022). We now discuss some ways to efficiently evaluate the $\Phi_{\lambda}$ and $\Phi_{\lambda}^{\pm}$-criteria, and hence their corresponding summarized versions. For a support size of $k$ and design X, one full evaluation of $\Phi_{\lambda}$ and $\Phi_{\lambda}^{\pm}$ requires consideration of all $\binom{p}{k}$ supports of size $k$. For $\Phi_{\lambda}^{\pm}$, an additional $2^{k}$ computations are required for each $\mathcal{A}$ to consider all sign vectors. Evaluating either criteria would benefit from a reduction in the number of supports considered. Sampling from $\mathcal{A}_{k}$ completely at random is intuitive, but may require a large number of samples to be representative. As high correlations between two factors can result in diminished lasso support recovery performance due to low $P(I_{\lambda}\,|\,\textbf{F},\,\boldsymbol{z})$, a support sampling method that balances (or approximately balances) the number of times pairwise sets of factors are included in $\mathcal{A}$ together can be advantageous. Smucker and Drew (2015) utilized nearly balanced incomplete block designs (NBIBDs) to approximate a model space with only a relatively small number of blocks, in the context of model-robust optimal designs. Hence, we recommend implementing the NBIBD sampling method to adequately represent $\mathcal{A}_{k}$ with between 64-128 supports for modest $p$ and $k$ values. We denote such a subset of supports by $\tilde{\mathcal{A}}_{k}$.
From Lemma 1, for a fixed $\mathcal{A}$, the probabilities are equal between $\boldsymbol{z}_{\mathcal{A}}$ and $-\boldsymbol{z}_{\mathcal{A}}$. Thus, only $2^{k-1}$ sign vectors need be considered for the $\Phi_{\lambda}^{\pm}$-criterion. While this cuts the computation in half, prior knowledge from the practitioner can be leveraged to select an even smaller set of representative sign vectors to evaluate. For example, if there is no knowledge on sign direction, the most likely sign vectors are those with an equal, or nearly equal, number of ±$\pm 1$. For $k=10$, one would need to only consider, up to reflection, the $126$ possible sign vectors having five $+1$’s and $-1$’s rather than all $5122^{10-1}=$ 512 vectors. There may also be strong prior belief about the signs of some or all of the factors that can reduce the set of sign vectors to a manageable number, or even to a single element.
@@ 5 Examples
We now demonstrate near-optimal SSDs found under our criteria via HILS under three example scenarios. For each case, the coordinate exchange algorithm was executed with $ntry=50$ random starting designs and identified $ncand$ final designs on the approximate Pareto front of the criteria described in Section 4. Of these $ncand$ designs, the final HILS design was selected using the $\Phi_{\text{max}}^{\pm}$ criterion. For each scenario, we provide the computational time that it took to find the HILS-optimal design on a Desktop with an Intel ® Core i7-6700 CPU @ 3.40GHz processor with 16GB RAM. We then compare the best identified designs by inspecting the $\Phi_{\lambda}$ and $\Phi_{\lambda}^{\pm}$ curves as a function of $\log(\lambda)$. These curves represent the average probability of sign recovery of the lasso solution path over a range of tuning parameter values. In comparing designs in this way, we eliminate the need to choose a tuning parameter. The design with larger $\Phi_{\lambda}$ or $\Phi_{\lambda}^{\pm}$ over a larger range of $\log(\lambda)$ is considered to be the best performing.
Additionally, we evaluate the performance of the HILS designs compared to competitors under $k$ and $\beta$ settings that are different to the $k$ and $\beta$ values used to construct the HILS designs. This evaluates the sensitivity of the HILS construction method to $k$ and $\beta$.
@@ 5.1 Scenario 1: $n=9$, $p=10$ Unknown Signs
The HILS algorithm was performed for $k=3$ and $\beta=3$ assuming unknown signs. The HILS-optimal design was selected in approximately six minutes of computation time. Because $n\approx p$, this scenario is close to one in which an orthogonal design exists. Therefore, this represents an experiment in which other optimal design criteria should perform rather well. For unknown signs, the optimal $c$ value is taken as $c^{*}=0$ following Theorem 3. Since $\textbf{C}^{*}$ is orthogonal and the $UE(s^{2})$ criterion targets orthogonality directly, the $UE(s^{2})$-optimal designs were deemed appropriate for comparison. Additionally, since the two criteria used to construct the DCDs are optimized under an orthogonal design, the DCD was also compared.
Figure 4 compares the $\Phi_{\lambda}^{\pm}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD (Singh and Stufken, 2022), and a $UE(s^{2})$-optimal design under a variety of $k$ and $\beta$ settings. The top-middle plot represents the $k$ and $\beta$ setting in which the HILS design was optimized for. The top row of Figure 4, representing $k=3$, demonstrates that the HILS design out performs the others in terms of lasso sign recovery regardless of the $\beta$ value. When $k=5$, Figure 4 shows little difference between the 3 designs. The HILS and $UE(s^{2})$- optimal designs show a larger $\Phi_{\lambda}^{\pm}$ for larger $\log(\lambda)$, but the DCD attains a slightly larger maximum of $\Phi_{\lambda}^{\pm}$ (for $\beta=3$ and $\beta=4$). Overall, the HILS-optimal design saw better performance than its competitors under the $k$ value that it was optimized for and comparable performance as the $k$ assumption was violated.
Figure 4: $\Phi_{\lambda}^{\pm}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD, and a $UE(s^{2})$-optimal design under a variety of $k$ and $\beta$ settings for scenario 1.

### v3

@@ 3.2 Optimal $\textbf{C}\in\mathcal{C}_{p,k}$ for relaxed criteria
Events (7) and (8) will be denoted here by $P(S_{\lambda}\,|\,c,\boldsymbol{\beta}_{\mathcal{A}})$ and $P(I_{\lambda}\,|\,c,\boldsymbol{z})$. The probabilities for the two events for $\textbf{C}\in\mathcal{C}_{p,k}$ are given in the following lemma:
@@
$\boldsymbol{u}$ $\sim N\left(\frac{\lambda}{1-c}\left[\boldsymbol{1}-z_{\mathcal{A}}\gamma\boldsymbol{z}_{\mathcal{A}}\right],\ \frac{1}{1-c}\left[\textbf{I}_{k}-\gamma\boldsymbol{z}_{\mathcal{A}}\boldsymbol{z}_{\mathcal{A}}^{T}\right]\right)\ ,$
$\boldsymbol{v}$ $\sim N\left(\lambda z_{\mathcal{A}}\gamma\boldsymbol{1},(1-c)\left[\textbf{I}+\gamma\textbf{J}\right]\right)\ ,$
@@
When $\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1}$ the random vectors $\boldsymbol{u}$ and $\boldsymbol{v}$ are distributed as
$\boldsymbol{u}\sim N\left(\frac{\lambda}{1+c(k-1)}\boldsymbol{1},\frac{1}{1-c}\left[\textbf{I}-\gamma\textbf{J}\right]\right),\qquad\boldsymbol{v}\sim N\left(\lambda k\gamma\boldsymbol{1},(1-c)\left[\textbf{I}+\gamma\textbf{J}\right]\right)\ ,$
for all $\mathcal{A}\in\mathcal{A}_{k}$. The corresponding criterion defined with respect to $\mathcal{C}_{p,k}$ and for a fixed $\lambda$ is denoted
$\psi_{\lambda}(c\,|\,k,\beta)=P(S_{\lambda}\,|\,c,\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1})\times P(I_{\lambda}\,|\,c,\boldsymbol{z})\ .$ (18)
@@
$\psi_{\lambda}^{\pm}(c\,|\,k,\beta)=\frac{1}{2^{k-1}}\sum_{\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}^{\pm}}P(S_{\lambda}\,|\,c,\boldsymbol{\beta}_{\mathcal{A}}=\beta\tilde{\boldsymbol{z}})\times P(I_{\lambda}\,|\,c,\boldsymbol{z}=\tilde{\boldsymbol{z}})\ .$ (19)
@@
The criteria $\psi_{\lambda}$ and $\psi_{\lambda}^{\pm}$, as well as their summarized versions, involve a single variable but are still challenging to optimize analytically. Fortunately, numerical optimization of these criteria is straightforward and computationally efficient. We demonstrate numerical optimization of these new criteria for the situation of $p=10$, $k=4$, and $\beta=2\sqrt{10}$. Figure 5 shows the contour plots of $\psi_{\lambda}$ and $\psi_{\lambda}^{\pm}$. The optimal $c$ values for the corresponding integral summary measures, $\psi_{\Lambda}$ and $\psi_{\Lambda}^{\pm}$, are $0.14$ and 00, respectively. The resulting optimal C matrices then match the ideal forms for the $\text{Var}(s+)$ and $\text{UE}(s^{2})$-criterion. While these ideal C are not possible for SSDs, this example provides further justification for these two heuristic SSD criteria for sign recovery under the lasso.
Figure 5: Contour plots of sign recovery probabilities when C is completely symmetric across possible values of the off-diagonal $c$ and $\log(\lambda)$ where $k=4$ factors are active with $\beta=2\sqrt{10}$. The red lines correspond to values of $c=0.14$ for $\psi_{\Lambda}$ and $c=0$ for $\psi_{\Lambda}^{\pm}$.
@@
We now study the behavior of $\psi_{\lambda}$ and $\psi^{\pm}_{\lambda}$ in a neighborhood about $c=0$ to address the potential optimality of orthogonal designs. Direct application of the multivariate Leibniz integral rule gives the following lemma.
@@
For all $\mathcal{A}$ where $|\mathcal{A}|=k$ and $\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1}>\boldsymbol{0}$, $\frac{d}{dc}P(I_{\lambda}|c,\boldsymbol{z})\big{|}_{c=0}=0$ for all $\lambda$ and $\frac{d}{dc}P(S_{\lambda}|c,\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1})\big{|}_{c=0}\geq 0$ for $\lambda$ satisfying
$2\lambda$ $\geq\frac{g(\tau)}{G(\tau)}\ ,$ (20)
where $g(\cdot)$ and $G(\cdot)$ refer to the standard Normal probability density function and cumulative distribution function, respectively, and $\tau=\beta-\lambda$.
A direct consequence of Lemma 3 is that $\psi_{\lambda}(0|k,\beta)$ can often be improved upon by some $c>0$. We state this important result as a theorem.
@@
Applying this result to $\psi_{\lambda}(c\,|\,k,\beta)$ shown in Figure 5, inequality (20) holds for $\log(\lambda)\geq-22.763$, which covers the entire region of $\lambda$ values for which $\psi_{\lambda}(c\,|\,k,\beta)>0$. It follows then that some $c>0$ will maximize the summary measures $\psi_{\Lambda}(c\,|\,k,\beta)$ or $\psi_{\max}(c\,|\,k,\beta)$. Theorem 2 thus provides a mathematically rigorous justification that the ideal designs sought after by the $\text{Var}(s+)$-criterion maximize the probability of sign recovery for $\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1}$. The investigations in Section 3.1 further suggest such designs may be superior when $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$, so long as the elements of $\boldsymbol{\beta}_{\mathcal{A}}$ are sufficiently large.
@@
$\begin{split}\frac{q}{\binom{k}{2}}\ \frac{\lambda g(\lambda)}{1-2G(-\lambda)}&\left(k(1-\lambda^{2})+(q-1)\frac{\lambda g(\lambda)}{1-2G(-\lambda)}\right)\leq\\ &\frac{g(\tau)}{G(\tau)}\left(\beta+\lambda+\lambda^{2}\tau-\left[\frac{\beta^{2}-\lambda^{2}}{2}+\beta\lambda\right]\frac{g(\tau)}{G(\tau)}\right)\end{split}$ (21)
where $q=p-k$.
Applying this result to $\psi_{\lambda}^{\pm}(c\,|\,k,\beta)$ shown in Figure 5, inequality (21) holds for $\log(\lambda)\in[-0.988,0.640]$. For $\log(\lambda)$ outside this region, there are some $c>0$ for which $\psi_{\lambda}^{\pm}(c)>0$, although the probabilities are relatively small. These small probabilities do not influence $\psi_{\max}^{\pm}$ and have minimal influence on $\psi_{\Lambda}^{\pm}$, making $c=0$ a likely global maximum for both criteria. We conjecture that generally $c^{*}=0$ and some $c^{*}>0$ are global maxima for $\psi_{\lambda}^{\pm}(c\,|\,k,\beta)$ and $\psi_{\lambda}(c\,|\,k,\beta)$, respectively.
@@ 4 Fast Design Construction with HILS
tr(V) and $\text{tr}(\textbf{V})\quad\text{ and }\quad||\textbf{C}-\textbf{C}^{*}||_{F}\ ,$
@@
The proposed construction algorithm first identifies a set of $ncand$ candidate designs through the Pareto front optimization search algorithm from Cao et al. (2017) based on the two lasso criteria. We then employ the exact lasso sign recovery criteria from Section 2 as a secondary criterion, as it more accurately measures a design’s screening capabilities, to rank the $ncand$ designs and choose the best one. We refer to this algorithm as the Heuristic-Initiated Lasso Sieve, or HILS.
@@
Although HILS can significantly reduce the number of evaluations of the relaxed lasso criteria compared to constructing an optimal design using, for instance, a coordinate exchange algorithm, there are settings when even a more modest number of evaluations is not computationally feasible. The same issue was encountered by Singh and Stufken (2023). We now discuss some ways to efficiently evaluate the $\Phi_{\lambda}$ and $\Phi_{\lambda}^{\pm}$-criteria, and hence their corresponding summarized versions. For a support size of $k$ and design X, one full evaluation of $\Phi_{\lambda}$ and $\Phi_{\lambda}^{\pm}$ requires consideration of all $\binom{p}{k}$ supports of size $k$. For $\Phi_{\lambda}^{\pm}$, an additional $2^{k-1}$ computations are required for each $\mathcal{A}$ to consider all sign vectors. Evaluating either criteria would benefit from a reduction in the number of supports considered. Sampling from $\mathcal{A}_{k}$ completely at random is intuitive, but may require a large number of samples to be representative. As high correlations between two factors can result in diminished lasso support recovery performance due to low $P(I_{\lambda}\,|\,\textbf{C},\,\boldsymbol{z})$, a support sampling method that balances (or approximately balances) the number of times pairwise sets of factors are included in $\mathcal{A}$ together can be advantageous. Smucker and Drew (2015) utilized nearly balanced incomplete block designs (NBIBDs) to approximate a model space with only a relatively small number of blocks, in the context of model-robust optimal designs. Hence, we recommend implementing the NBIBD sampling method to adequately represent $\mathcal{A}_{k}$ with between 64-128 supports for modest $p$ and $k$ values.
From Lemma 1, for a fixed $\mathcal{A}$, the probabilities are equal between $\boldsymbol{z}_{\mathcal{A}}$ and $-\boldsymbol{z}_{\mathcal{A}}$. Thus, only $2^{k-1}$ sign vectors need be considered for the $\Phi_{\lambda}^{\pm}$-criterion. While this cuts the computation in half, prior knowledge from the practitioner can be leveraged to select an even smaller set of representative sign vectors to evaluate. For example, if there is no knowledge on sign direction, the most likely sign vectors are those with an equal, or nearly equal, number of ±$\pm 1$. For $k=10$, one could reduce the number of sign vectors from $5122^{9}=$ 512 to only the $126$ possible sign vectors having five $+1$’s and $-1$’s. There may also be strong prior belief about the signs of some or all of the factors that can reduce the set of sign vectors to a manageable number, or even to a single element.
@@ 5 Examples
We now demonstrate near-optimal SSDs found under the criteria via HILS under three example scenarios. For each case, the coordinate exchange algorithm was executed with $ntry=50$ random starting designs and identified $ncand$ final designs on the approximate Pareto front of the criteria described in Section 4. Of these $ncand$ designs, the final HILS design was selected using the $\Phi_{\text{max}}^{\pm}$ criterion. For each scenario, we provide the computational time that it took to find the HILS-optimal design on a Desktop with an Intel ® Core i7-6700 CPU @ 3.40GHz processor with 16GB RAM. Additionally, we evaluate the performance of the HILS designs compared to competitors under $k$ and $\beta$ settings that are different to the $k$ and $\beta$ values used to construct the HILS designs. This evaluates the sensitivity of the HILS construction method to $k$ and $\beta$.
[…]
@@ 5.1 Scenario 1: $n=9$, $p=10$ Unknown Signs
The HILS algorithm was performed for $k=3$ and $\beta=3$ assuming unknown signs. The HILS-optimal design was selected in approximately six minutes of computation time. Because $n\approx p$, this scenario is close to one in which an orthogonal design exists. Therefore, this represents an experiment in which optimal designs under traditional SSD criteria should perform rather well. For unknown signs, we set $c^{*}=0$ following Theorem 3. Since $\textbf{C}^{*}$ is orthogonal and the $\text{UE}(s^{2})$ criterion targets orthogonality directly, the $\text{UE}(s^{2})$-optimal designs were deemed appropriate for comparison. Additionally, since the two criteria used to construct the DCDs are optimized under an orthogonal design, the DCD was also compared.
Figure 7 compares the $\Phi_{\lambda}^{\pm}$ curves as a function of $\log(\lambda)$ for the hypothetically optimal completely symmetric (CS optimal) design, HILS-optimal design, DCD, and a $\text{UE}(s^{2})$-optimal design under a variety of $k$ and $\beta$ settings. The top-middle plot represents the $k$ and $\beta$ setting in which the HILS design was optimized for. The top row of Figure 7, representing $k=3$, demonstrates that the HILS design out performs the others in terms of lasso sign recovery regardless of the $\beta$ value. When $k=5$, Figure 7 shows little difference between the 3 designs. The HILS and $\text{UE}(s^{2})$- optimal designs show a larger $\Phi_{\lambda}^{\pm}$ for larger $\log(\lambda)$, but the DCD attains a slightly larger maximum of $\Phi_{\lambda}^{\pm}$ (for $\beta=3$ and $\beta=4$). Overall, the HILS-optimal design saw better performance than its competitors under the $k$ value that it was optimized for and comparable performance as the $k$ assumption was violated.
Figure 7: $\Phi_{\lambda}^{\pm}$ curves as a function of $\log(\lambda)$ for the HILS optimal design, DCD, and a $\text{UE}(s^{2})$-optimal design under a variety of $k$ and $\beta$ settings for scenario 1. The HILS optimal design shown in all panels was chosen assuming $k=3$ and $\beta=3$. The CS Optimal curve refers to the hypothetical design having the optimal completely symmetric form for $k=3$ and $\beta=3$.